OpenStudy (anonymous):

Problem 3.17 A tiger leaps horizontally from a 6.4m -high rock with a speed of 4.0m/s . How far from the base of the rock will she land? Express your answer using two significant figures.

OpenStudy (aaronq):

 first find how long it takes for the tiger to hit the ground using the acceleration due to gravity. Next use the time and the horizontal velocity to find the distance travelled. You'll need some of these: \(\sf v_f=v_i+a\Delta t\) \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf v_f^2=v_i^2+2a_x\Delta x\)